Solve 2-143 please
(2-7 Bayes Theorem) Suppose that P(AIB) 0.6, P() 0.4, and P(B 0.3 Determine P(BIA). 2-143. Suppose that P(AIB)-0.5, P(AIB) 0.1, and P(B)- 0.7. Determine P(BIA).
Answer #1

Solution:

Given in the question

P(B/A) = P(A intersection B)/P(A)

P(A/B) = P(A intersection B)/P(B)

P(B/A) =P(A/B)*P(B)/P(A)

P(A) =P(A/B)*P(B)/P(B/A) = 0.5*0.7/P(B/A) = 0.35/P(B/A)

P(A/B') = P(A)((1-P(B/A))/((1-P(B))

0.1*P(B/A)=0.35(1-P(B/A))/(1-0.7)

0.03P(B/A) = 0.35-0.35P(B/A)

0.38P(B/A) =0.35

P(B/A) =0.35/0.38 = 0.92